# Cosmology with non-conformal holographic matter

### v2

@@ 1 Introduction
Ordinary matter described by the Standard Model is believed to comprise only a small fraction of the total energy density in the universe. The remaining components, dubbed dark matter and dark energy, are still mysterious hawking_ellis_1973 ; weinberg_cosmology_2008 ; Frieman2008 ; 1940oort . In the $\Lambda$CDM model of cosmology, the dark matter is assumed to be pressureless matter with equation of state parameter $w=0$, while the dark energy is taken to be a cosmological constant with $w=-1$ Bahcall_1999 . [^1] The equation of state parameter is defined as $w=p/e$ where $p$ is the pressure and $e$ is the energy density. In general, this can depend on the temperature. But the dark sector might well be more complicated or include additional components. It is common in observational studies to consider phenomenological models with different, possibly time-dependent equation of state parameters and to constrain the possible behaviors of these darkenergy ; dark_energy .
@@
There have been many other works in the past investigating the thermodynamic behavior of holographic field theories or aspects of their cosmological physics. Papers that specifically consider aspects of equilibrium and non-equilibrium thermodynamics for non-conformal holographic theories include Attems_2016 ; G_rsoy_2018 ; Kleinertt_2016 ; Attems_2017 ; Ballon_QCD_2021 ; Buchel_QCD_2016 ; G_rsoy_QCD_2016 . Various papers have studied the behavior of holographic field theories on fixed cosmological backgrounds, including Ghosh_2018 ; Casalderrey_Solana_2021 ; Marolf_2011 . Other papers that have studied cosmological evolution of holographic field theories dynamically coupled to gravity include 2022paper ; dynamicalinflaton ; Koyama2001 where the authors go beyond the quasi-static approximation to study situations where the quantum field theory may be out of equilibrium. Holographic non-conformal matter has also been studied previously in the context of braneworld cosmology, with equation-of-state parameters arising from scalar potentials different from those considered in this paper (for related work, see Fichet:2022ixi ; Fichet:2022xol ; Fichet:2023dju ). In reheating ; reheating2 ; reheating3 ; warminflation the authors consider holographic field theories as part of the physics of inflation; an important aspect of strongly coupled QFTs is that they quickly thermalize and enter the hydyodynamic regime making them potentially useful for facilitating the exit from inflation to the hot Big Bang. Novel aspects of our paper compared to previous works include i) a detailed investigation of how the scale dependent equation of state parameter depends on parameters (dimension of perturbing operator and length of RG flow) in perhaps the simplest possible example of a holographic field theory with an RG flow between UV and IR CFT ii) a different numerical method that does not make use of the counterterms in the gravitational action and iii) a numerical study of the dynamical evolution of the scale factor for flat cosmologies where the only matter is the non-conformal holographic theory.
@@ 2.1 Equilibrium thermodynamics in Minkowski space
where $\textit{S}_{0}$ is the CFT action and J is a spacetime-independent source. In order to study the thermodynamics, we consider the Euclidean theory on $\mathbb{R}^{3}\times S^{1}$, where the thermal circle has length $\beta=1/T$. From dimensional analysis, $[J]=4-\Delta$, so we can define a dimensionless temperature parameter
@@
In the holographic field theories that we consider below, it will be simplest to read off the entropy density as a function of temperature from the gravity solutions, so it will be convenient to express the other thermodynamic quantities in terms of this. Starting from the First Law of Thermodynamics $dE=TdS-pdV$ and the definition of free energy $F=E-TS$, we have that $dF=-SdT-pdV$, so that
@@
[…]
@@ 2.2 Cosmology
For conformal matter, recalling that ${\cal S}$, ${\cal E}$, and ${\cal P}$ are constant, we have from equation (15) that
@@
$-{\ddot{a}\over a}={4\pi G\over 3}(3p+e)$ (24)
@@
where $w=p/e={\cal P}/{\cal E}$ so the deviations of scale factor deceleration from the CFT behavior $q=-1$ are directly related to the deviations of the equation of state parameter $w$ from the CFT behavior $w=1/3$. Calculating $w$ as a function of ${\cal T}$ for various holographic field theories will be one of our main goals below.
@@ 3 Holographic calculations
where we have set $L_{AdS}=1$, giving $\Lambda=-6$ for the $4+1$ dimensional asymptotically AdS background. We also have the scalar field equation of motion $-\nabla^{2}\phi+V^{\prime}(\phi)=0$ which leads to
@@ 5 Discussion
Using the framework of AdS/CFT, we have characterized the effects on cosmological evolution of strongly coupled non-conformal quantum field with an RG flow between UV and IR CFTs. The analysis showed a very particular signature: the equation of state parameter approaches the conformal value $w=1/3$ for large and small temperatures but exhibits a dip to lower values in the regime where temperatures are similar to the scale set by the source for the relevant operator. While there is no particular reason to expect that such a strongly coupled non-conformal theory is present in the dark sector, it is interesting to keep this in mind as a possibility since it gives a behavior more general than the more standard matter, radiation or cosmological constant.
Our analysis was restricted to the quasi-static regime in which equilibrium physics is assumed. We have provided evidence that this approximation is reasonable for the theories we have considered. For more accuracy, or in cases where (perhaps because of other matter components) the cosmological evolution becomes fast relative to the time scale associated with the field theory, methods similar to 2022paper ; dynamicalinflaton ; Koyama2001 could be used in order to study the evolution according to the full semiclassical Einstein equation.

### v3

@@ 1 Introduction
[…]
@@
There have been many other works in the past investigating the thermodynamic behavior of holographic field theories or aspects of their cosmological physics. Papers that specifically consider aspects of equilibrium and non-equilibrium thermodynamics for non-conformal holographic theories include Attems_2016 ; G_rsoy_2018 ; Kleinertt_2016 ; Attems_2017 ; Ballon_QCD_2021 ; Buchel_QCD_2016 ; G_rsoy_QCD_2016 . Various papers have studied the behavior of holographic field theories on fixed cosmological backgrounds, including Ghosh_2018 ; Casalderrey_Solana_2021 ; Marolf_2011 . Other papers that have studied cosmological evolution of holographic field theories dynamically coupled to gravity include 2022paper ; dynamicalinflaton ; Koyama2001 where the authors go beyond the quasi-static approximation to study situations where the quantum field theory may be out of equilibrium. Holographic non-conformal matter has also been studied previously in the context of braneworld cosmology, with equation-of-state parameters arising from scalar potentials different from those considered in this paper (for related work, see Fichet:2022ixi ; Fichet:2022xol ; Fichet:2023dju ). In reheating ; reheating2 ; reheating3 ; warminflation the authors consider holographic field theories as part of the physics of inflation; an important aspect of strongly coupled QFTs is that they quickly thermalize and enter the hydyodynamic regime making them potentially useful for facilitating the exit from inflation to the hot Big Bang.
Novel aspects of this work include i) a systematic study of how the scale dependent equation of state parameter varies with the scaling dimension of the perturbing operator and with the length of the RG flow, using perhaps the simplest holographic QFT that interpolates between UV and IR CFTs; ii) the use of a numerical method that obtains thermodynamic quantities directly from bulk solutions without introducing explicit holographic counterterms. While this technique has appeared before in other contexts (see, e.g. G_rsoy_2009 ), here we adapt it to track cosmological solutions along the RG flow and to map out the full $w(\mathcal{T})$ curve; and iii) a numerical exploration of the dynamical evolution of the scale factor in spatially flat cosmologies where the only matter component is the strongly coupled non-conformal holographic theory.
@@ 2.1 Equilibrium thermodynamics in Minkowski space
where $\textit{S}_{0}$ is the CFT action and J is a spacetime-independent source. In order to study the thermodynamics, we consider the Euclidean theory on $\mathbb{R}^{3}\times S^{1}$, where the thermal circle has length $\beta=1/T$. From dimensional analysis, $J]=4-\Delta$, so we can define a dimensionless temperature parameter
@@
In the holographic field theories that we consider below, it will be simplest to read off the entropy density as a function of temperature from the gravity solutions, so it will be convenient to express the other thermodynamic quantities in terms of this. Starting from the First Law of Thermodynamics dE=TdSpdVdE=TdS-pdV and the definition of free energy F=ETSF=E-TS, we have that dF=SdTpdVdF=-SdT-pdV, so that
@@
$\mathcal{P}(\mathcal{T})=1/\mathcal{T}^{4}\int_{0}^{\mathcal{T}}\tilde{\mathcal{T}}^{3}\mathcal{S}(\tilde{\mathcal{T}})d\tilde{\mathcal{T}}\qquad\qquad\mathcal{E}(\mathcal{T})=\mathcal{S}(\mathcal{T})-\mathcal{P}(\mathcal{T})$ (8)
@@ 2.2 Cosmology
[…]
@@
${\ddot{a}\over a}={4\pi G\over 3}(3p+e)$ (24)
@@
where $w=p/e={\cal P}/{\cal E}$ so the deviations of scale factor deceleration from the CFT behavior q=1q=-1 are directly related to the deviations of the equation of state parameter $w$ from the CFT behavior $w=1/3$. Calculating $w$ as a function of $\cal T$} for various holographic field theories will be one of our main goals below.
@@ 3 Holographic calculations
where we have set $L_{AdS}=1$, giving $\Lambda=-6$ for the $4+1$ dimensional asymptotically AdS background. We also have the scalar field equation of motion $\nabla^{2}\phi+V^{\prime}(\phi)=0$ which leads to
@@ 4.1 Validity of the quasi-static approximation
4.2 Trace–anomaly effects in the quasi–static regime
For a CFT (or a QFT that behaves as a CFT in the UV) on an FRW background, there are additional contributions to the stress-energy tensor due to the conformal anomaly. We have ignored these so far, since we expect them only to be relevant at very early times before the relevant perturbation comes into play.
In this subsection we analyze more explicitly when this is justified, finding that as long as
$J^{\frac{1}{4-\Delta}}\;\ll\;\frac{M_{\rm Pl}}{N}\;,$ (45)
the transient anomaly–driven effects always die out before the relevant deformation becomes important.
Holographic stress tensor and trace anomaly.
Using the standard AdS5/CFT4 dictionary $l^{3}/G_{5}=2N^{2}/\pi$, the renormalised boundary stress tensor on a spatially flat FRW metric $\mathrm{d}s^{2}=-\mathrm{d}t^{2}+a^{2}(t)\,\mathrm{d}\vec{x}^{2}$ is
[…]
where $E\geq 0$ is an integration constant fixed by initial conditions. Taking the trace of (46) reproduces the CFT4 anomaly Henningson_1998 ; Balasubramanian_1999
$\langle T^{\mu}_{\;\mu}\rangle=-\frac{N^{2}}{\pi^{2}}\!\left(E_{4}+I_{4}\right),$ (47)
with $E_{4}=R_{\mu\nu\rho\sigma}R^{\mu\nu\rho\sigma}-4R_{\mu\nu}R^{\mu\nu}+R^{2}$ (the four–dimensional Euler density) and $I_{4}=C_{\mu\nu\rho\sigma}C^{\mu\nu\rho\sigma}$ (the square of the Weyl tensor). For a flat FRW background $I_{4}=0$, and (47) reduces to $\langle T^{\mu}_{\;\mu}\rangle=-\frac{3N^{2}}{16\pi^{2}}\,H^{2}(\dot{H}+H^{2}),$ where $H=\dot{a}/a$.
Einstein equation and exact time dependence.
Coupling (46) to four-dimensional gravity via $G^{0}_{\;0}=8\pi G_{N}T^{0}_{\;0}$ gives
$a^{2}=\frac{N^{2}G_{N}}{4\pi}\Bigl{(}\dot{a}^{2}+E/\dot{a}^{2}\Bigr{)}.$ (48)
Solutions with different values of $E$ are related to each other by a rescaling $a\to E^{1/4}a$), so we will choose a value $E_{\star}=4\pi^{2}/(G_{N}^{2}N^{4})\;$. The equation (48) has two branches of solutions, an inflating branch and one associated with ordinary late-time cosmological evolution. We are interested in the latter, which is governed by
$\dot{a}^{2}=\frac{2\pi}{G_{N}N^{2}}\Bigl{(}a^{2}-\sqrt{a^{4}-1}\Bigr{)}.$ (49)
Early and late times.
A characteristic feature of the solution is that there is some initial time $t_{0}$ with finite energy density for which the solution cannot be extended to earlier times.
Our choice of $E_{*}$ corresponds to taking $a=1$ at this initial time. We have energy density $T^{0}_{\;0}(t_{0})=-3/(4N^{2}G_{N}^{2})$ here; this is the maximum energy density along the trajectory. For $a\gg 1$ the solution expands as
$\dot{a}^{2}\simeq\frac{\pi}{G_{N}N^{2}\,a^{2}}\quad\Longrightarrow\quad a^{2}(t)\simeq 2\sqrt{\frac{\pi}{G_{N}N^{2}}}\;t,$
so the universe asymptotes to the standard radiation law $a\propto t^{1/2}$.
Decay of the anomaly contribution.
Splitting $T^{0}_{\;0}$ into an “anomaly” part $\rho_{A}\propto\dot{a}^{4}/a^{4}$ and a “radiation” part $\rho_{R}\propto E/a^{4}$, equation (49) implies
$\frac{\rho_{A}}{\rho_{R}}=\frac{\dot{a}^{4}}{E}=\Bigl{(}a^{2}-\sqrt{a^{4}-1}\Bigr{)}^{2}\xrightarrow{a\gg 1}\frac{1}{4\,a^{4}}.$
Hence the anomaly contribution is approximately equal to the radiation term at the initial time and is suppressed by an extra factor $a^{-4}$ at later times. Taking the equality point as a marker, the energy density at the end of the anomaly era is
$\rho_{\text{end}}\;\simeq\;\epsilon\frac{1}{G_{N}^{2}N^{2}},$ (50)
for some small $\epsilon$.
Relevant deformation versus anomaly.
For a relevant deformation $J\mathcal{O}$ with $\cal O$} of dimension $\Delta<4$ we have an associated energy scale $\rho_{J}\sim N^{2}J^{\,4/(4-\Delta)}$. To ensure that the RG flow becomes important only after the trace–anomaly can be ignored, we require
$\rho_{J}\;<;\rho_{\text{end}}\;\simeq\;\epsilon\frac{1}{G_{N}^{2}N^{2}}\quad\Longrightarrow\quad J^{\frac{1}{4-\Delta}}\;\ll\;\frac{M_{\text{pl}}}{N}.$
Thus, provided that the energy scale associated with relevant perturbation is below the Planck scale by at least a factor of $N$, we can ignore the effects of the conformal anomaly for our analysis.
@@ 5 Discussion
Using the framework of AdS/CFT, we have characterized the effects on cosmological evolution of a strongly coupled non-conformal quantum field theory whose RG flow interpolates between UV and IR CFTs. The analysis showed a distinctive signature: the equation–of–state parameter approaches the conformal value $w=1/3$ at very high and very low temperatures, but dips to smaller values when the temperature is comparable to the scale set by the source for the relevant operator.
The temporary dip has a simple physical origin. As the system cools to the scale associated with the relevant perturbation, some of the degrees of freedom of the CFT begin to behave more like massive non-relativistic matter with $w=0$ before dropping out completely in the IR. Once the temperature falls far enough that the theory is governed by the IR fixed point, the remaining excitations are again effectively massless, radiation dominates, and $w$ returns to $1/3$. The depth of the dip is governed by how drastically the number of active relativistic degrees of freedom is reduced along the RG flow, and this is captured by the change of the bulk potential $V(\phi_{\min})$. For fixed $c$, decreasing the scaling dimension $\Delta$ makes the deformation more relevant and therefore eliminates a larger fraction of UV degrees of freedom, pushing the minimum of $w$ lower. For fixed $\Delta$, decreasing $c$ has a similar effect: it deepens the potential and therefore gives a larger fraction of the degrees of freedom that are removed during the RG-flow between the UV and IR theories, again producing a more pronounced reduction in $w$ in the intermediate stages.
While there is no particular reason to expect that such a strongly coupled non-conformal theory exists in the dark sector, it is useful to keep this possibility in mind since it gives a behavior more general than the standard matter, radiation or cosmological constant.
Our analysis was restricted to the quasi-static regime in which equilibrium physics is assumed. We have provided evidence that this approximation is reasonable for the theories we have considered. We also examined the effects of the trace anomaly and argued that we can ignore this during the period of evolution where the RG-flow is relevant provided that $J^{\frac{1}{4-\Delta}}\ll\frac{M_{\text{pl}}}{N}$. In this case, the conformal anomaly only affects the very early evolution where the energy density is high enough that the relevant perturbation is unimportant.
For more accuracy, or in cases where (perhaps because of other matter components) the cosmological evolution becomes fast relative to the time scale associated with the field theory, methods similar to 2022paper ; dynamicalinflaton ; Koyama2001 could be used to study the evolution using the full semiclassical Einstein equation.